OpenStudy (anonymous):

13=n/6+4 n=? please help

jhonyy9 (jhonyy9):

- so if the denominator is just 6 this will be 13= n/6 +4 13*6=n +4*6 78=n+24 54=n n=54

jhonyy9 (jhonyy9):

13=54/6 +4 13=9+4 13=13